# Dark Dwarfs: Dark Matter-Powered Sub-Stellar Objects Awaiting Discovery at the Galactic Center

### v1

@@ Abstract
${}^{7}\textrm{Li}+\textrm{p}$ $\rightarrow^{4}\textrm{He}+^{4}\textrm{He}.~{}$ (12)
Sub-stellar objects with masses lighter than $0.35{\rm M}_{\odot}$ are fully convective Chabrier and Baraffe (1997), implying that the ratios of lithium to hydrogen are constant throughout the star. The depletion rate is then given by Bildsten et al. (1997); Ushomirsky et al. (1998)
[…]
where $X$ is the hydrogen mass fraction, $m_{p}$ is the proton mass, $f$ is the lithium to hydrogen ratio, and $\sigma$ the relevant cross section, which we take from Caughlan and Fowler (1988). We can solve Eq. (13) using our analytic model but, because the nuclear burning rate (12) is highly sensitive to the core temperature, the model needs to be refined to include the corrections from Coulomb pressure in the stellar core. We explain our method for this in Appendix I.4.
In Fig. 5 we show the resulting survival fraction after 1 Gyr. In the case of the SM, we find that Li-7 is depleted in objects heavier than $0.062{\rm M}_{\odot}$, consistent with other theoretical predictions Pozio (1991); Nelson et al. (1993); Chabrier et al. (1996); Phillips et al. (2020). In contrast, a significant fraction of lithium-7 survives in dark dwarfs of this mass and heavier in regions where $\rho_{\rm DM}\gtrsim 10^{5}$ GeV/cm3. Therefore, given an observed mass and estimated age, the spectroscopic detection of lithium can serve as a marker for DM heating.
@@
Conclusion
We have found that the structure, evolution, and fate of sub-stellar objects in the presence of dark matter annihilation is markedly different than the Standard Model predictions. Objects lighter than the hydrogen burning limit begin their lives as brown dwarfs but ultimately evolve to become dark dwarfs — eternal objects powered by DM burning. Their radii, temperatures, and luminosity are constant. The H-burning limit itself is modified, depending on the incident DM flux. Stars heavier than this behave similarly to red dwarfs but are predicted to be larger and brighter than SM red dwarfs of identical mass due to the additional contribution from DM heating. Their core temperatures can be reduced, causing them to retain lithium in mass ranges where the SM predicts it is depleted. The detection of lithium-7 in objects heavier than the lithium burning limit would provide evidence for the existence of DM heating. The optimum location to search for these signatures is towards the galactic center, where low DM velocities and large densities maximize the incoming DM flux.
@@ Acknowledgements
Dark Dwarfs:
Dark Matter-Powered Sub-Stellar Objects Awaiting Discovery at the Galactic Center
Supplemental Material
Djuna Croon, Jeremy Sakstein, Juri Smirnov, Jack Streeter
I Analytic Model
In this section we review the derivation of the analytic model that we adopted for our calculations above, explore the effects of the assumptions we have made, and derive the range of validity of our model.
I.1 Polytrope Model for the Interior
We begin by reviewing the polytropic description of sub-stellar objects, following reference Auddy et al. (2016). The EOS for the combination of electron degeneracy pressure $P_{\rm deg}$ and the gas pressure of the ions $P_{\rm ion}$ is an $n=3/2$ polytrope:
$P=K\rho^{\frac{5}{3}}\textrm{ with }K=C\mu_{e}^{-5/3}(1+\gamma(\psi)+\alpha\psi).$ (S1)
@@ I.1 Polytrope Model for the Interior
$C=\frac{2}{5}aA^{\frac{5}{2}},~{}A=\frac{(3\pi^{2}\hbar^{3}N_{A})^{\frac{2}{3}}}{2m_{e}},~{}a=\frac{2}{3}\frac{4\pi m_{e}^{\frac{3}{2}}}{(2\pi\hbar)^{3}};$ (S2)
@@
$\gamma(\psi)=-\frac{5}{16}\psi\ln(1+e^{-\frac{1}{\psi}})+\frac{15}{8}\psi^{2}\left[\frac{\pi^{2}}{3}+\mathrm{Li}_{2}(-e^{-\frac{1}{\psi}})\right]$ (S4)
with $\mathrm{Li}_{2}$ the polylogarithm of order-2. Using the properties of $n=3/2$ polytropes, it is possible to find expressions for the radius, central temperature, central pressure, and central density as a function of the degeneracy parameter $\psi$ Auddy et al. (2016):
$R$ $=R_{0}\mu_{e}^{-5/3}(1+\gamma(\psi)+\alpha\psi)M_{0}^{-\frac{1}{3}}\,,$ (S5)
$T_{c}$ $=T_{0}M_{0}^{\frac{4}{3}}\mu_{e}^{\frac{8}{3}}\frac{\psi}{(1+\gamma(\psi)+\alpha\psi)^{2}}\,,$ (S6)
$\rho_{c}$ $=\rho_{0}M_{0}^{2}\frac{\mu_{e}^{5}}{(1+\gamma(\psi)+\alpha\psi)^{3}}\,,$ (S7)
$P_{c}$ $=P_{0}M_{0}^{\frac{10}{3}}\frac{\mu_{e}^{\frac{20}{3}}}{(1+\gamma(\psi)+\alpha\psi)^{4}}\,,$ (S8)
@@ I.2 Photosphere Models
Table S1: Parameter points for the phase transition Chabrier et al. (1992). Note that points G and H are very close to the second critical point of hydrogen, where a liquid-liquid phase transition occurs.
The photosphere, which defines the effective temperature and the stellar luminosity, is the point at which the optical depth $\tau$ falls to $2/3$. At this point, there is a phase transition from the mixture of degenerate electrons and ionic gas described above to a phase of molecular hydrogen. Modeling the latter, one can derive expressions for the effective temperature and luminosity. Several parameter points given in table S1 span the possible range of surface temperatures at which the phase transition takes place:
$T_{\rm eff}=b_{1}\times 10^{6}\rho_{e}^{2/5}\psi^{\nu}\rm K,$ (S9)
where $b_{1}$ and $\nu$ are parameters of the model Chabrier et al. (1992). Solving the condition $\tau=2/3$ yields an expression for the pressure at the photosphere, which can be used to find the density and effective temperature from the idea gas law (see Auddy et al. (2016) for the details – here we generalise the solution presented in this work). The result is:
$T_{\rm eff}=1.57466\times 10^{4}\,b_{1}\psi^{\nu}\left(\frac{M_{0}^{5/3}\psi^{-\nu}}{b_{1}\kappa_{R}(1+\gamma+\alpha\psi)^{2}}\right)^{2/7}\rm K,$ (S10)
@@
$\frac{L_{\rm surf}}{L_{\odot}}=\frac{0.0578353b_{1}^{3}M_{0}}{\kappa_{R}}\psi^{3\nu}\left(\frac{M_{0}^{5/3}\psi^{-\nu}}{b_{1}\kappa_{R}(1+\gamma+\alpha\psi)^{2}}\right)^{1/7}.$ (S11)
@@
I.3 Sensitivity to the Molecular Hydrogen Phase Transition
@@
Figure S1: Minimum mass for hydrogen burning (using the criterion $L_{\rm DM}=L_{\rm HB}$) as a function of DM ambient density for different models of the photosphere. Here we use $v_{\rm DM}=$50 km/s.
@@
Our calculation relies upon a choice of parameter point for the phase transition, which introduces a source of theoretical uncertainty into our calculations. We explore this in this section. A convenient measure for quantifying the model-dependency is the minimum mass for hydrogen burning, which is DM density and velocity dependent. Since DM annihilation reduces the core temperature and density below the threshold for PP-burning, the MMHB is increased whenever DM burning is important. Demanding $L_{\rm HB}=L_{\rm DM}$ leads to the minimum mass shown in figure S1 for the different parameter points given in table S1. The figure demonstrates that the model’s predictions only significantly diverge for very large DM densities (with the exception of edge cases G and H), implying that the choice of parameter point is not an important source of uncertainty for our conclusions.
I.4 Coulomb Corrections to the Equation of State
@@
At low temperatures, the effects of Coulomb repulsion becomes important and the EOS deviates from an $n=3/2$ polytrope. This has two consequences for our work. First, it limits the applicability of the analytic model we have assumed to high mass objects and, second, the exponential sensitivity of the lithium-7 burning rate in Eq. (13) mandates that we include these corrections to calculate the core temperature accurately. We derive the mass range where our results are valid and expressions for the corrected core temperature in this section.
The effects of Coulomb interactions depend upon the plasma parameter Clayton (1983)
$\Gamma=\frac{e^{2}}{k_{B}T}\left(\frac{4\pi\rho}{3m_{p}}\right)^{\frac{1}{3}},$ (S12)
@@
where $e$ is the electron charge and $m_{p}$ is the proton mass. Early in the star’s evolution when lithium is being depleted the plasma parameter $\Gamma\ll 1$, corresponding to the Deybe-Hückel regime Ushomirsky et al. (1998). In this limit, each ion can be thought of as being surrounded by a spherically symmetric but inhomogeneously charged cloud that screens its charge. The Coulomb pressure of these screened charges is Rebolo et al. (1992)
$P_{\rm C}=-\frac{e^{2}}{3}\left(\frac{\pi}{k_{B}T}\right)^{\frac{1}{2}}\left(\frac{\rho\zeta}{m_{p}}\right)^{\frac{3}{2}};~{}\textrm{with }\zeta=\sum_{j}\frac{X_{j}}{A_{j}}Z_{j}(1+Z_{j}),$ (S13)
@@
where $j$ runs over all ions with mass fraction $X_{j}$, atomic number $A_{j}$ and charge $Z_{j}$$\zeta=1.875$ for our benchmark model. The ratio of the Coulomb pressure to the central pressure calculated using the polytrope model for various DM densities is shown in figure S2. In the SM, the polytropic approximation ceases to be valid ($|P_{\rm C}|>P_{c}$) when $M<0.18{\rm M}_{\odot}$. This threshold is reduced when DM annihilation is present. The results we presented in this work correspond to heavier objects.
@@
To incorporate the effects of the Coulomb pressure into our lithium calculation, we use that fact that $T_{c}\propto\mu_{\rm eff}^{2}$ Bildsten et al. (1997); Ushomirsky et al. (1998) with $\mu_{\rm eff}=k_{B}N_{A}\rho_{c}T_{c}/P_{c}$. At fixed central density, the negative Coulomb pressure reduces the total pressure from the polytrope model’s prediction, which raises the central temperature. To incorporate this effect, we first calculate $T_{c}$ using the polytropic model and then scale this by a factor of $P_{c}^{2}/(P_{c}+P_{\rm C})^{2}$ to account for the correction to $\mu_{\rm eff}$.
@@
At late times, when the star has cooled and reached its equilibrium state, either red or dark dwarf, the plasma parameter has evolved to $\Gamma\gg 1$ and the Debye-Hückel approximation does not apply. Instead, the plasma is comprised of a sea of degenerate electrons surrounding a lattice of ions arranged to maximize their separation, minimizing their Coulomb repulsion Clayton (1983). In this limit, the Coulomb pressure can be found using the the Wigner-Seitz approximation where each ion with atomic number $Z$ sits at the center of a neutral sphere containing $Z$ electrons. The Coulomb pressure is Clayton (1983)
@@
$\begin{split}P_{\rm C}&=K_{\rm C}\rho^{\frac{4}{3}};\\ K_{\rm C}&=-\frac{3}{10}\left(\frac{4\pi\langle Z^{2}\rangle}{3}\right)^{\frac{1}{3}}e^{2}\left(\frac{N_{A}}{\mu_{e}}\right)^{\frac{4}{3}}\approx 5.72\times 10^{12}\textrm{g/cm/s}^{2}.\end{split}$ (S14)
@@
Comparing the central Coulomb pressure to the central pressure (S8), the latter exceeds the former in objects where $M\lesssim 10^{-3}\,{\rm M}_{\odot}\sim\rm M_{\rm jupiter}$, independent of the DM density. Thus, the equilibrium configurations we have studied are valid at even lower masses than estimated above, the caveat being that the polytrope model provides an invalid description of their formation.
@@
Figure S2: Ratio of the Coulomb to central pressure as a function of mass for varying DM densities. The polytrope model presented in this work becomes invalid when $|P_{\rm C}|/P_{c}>1$.

### v2

@@ 4 Lithium burning
$^{7}\textrm{Li}+\textrm{p}$ $\rightarrow^{4}\textrm{He}+^{4}\textrm{He}.$~ (4.1)
Objects with masses lighter than $0.35{\rm M}_{\odot}$ are fully convective [109], implying that the ratios of lithium to hydrogen are constant throughout the star. The depletion rate is then given by [110, 111]
$\frac{d\ln f}{dt}=-\frac{4\pi X}{Mm_{p}}\int_{0}^{R}\rho^{2}\langle\sigma v\rangle r^{2}dr$ (4.2)
where $X$ is the hydrogen mass fraction, $m_{p}$ is the proton mass, $f$ is the lithium to hydrogen ratio, and $\sigma$ the relevant cross section, which we take from [112]. We can solve Eq. (4.2) using our analytic model but, because the nuclear burning rate (4.1) is highly sensitive to the core temperature, the model needs to be refined to include the corrections from Coulomb pressure in the stellar core. We explain our method for this in Appendix A.4.
In Fig. 5 we show the resulting survival fraction after 1 Gyr; results for other benchmark ages can be computed using our accompanying code. In the case of the SM, we find that Li-7 is depleted in objects heavier than $0.062{\rm M}_{\odot}$, consistent with other theoretical predictions [113, 114, 115, 116]. In contrast, a significant fraction of lithium-7 survives in dark dwarfs of this mass and heavier in regions where $\rho_{\rm DM}\gtrsim 10^{5}$ GeV/cm3. Therefore, given an observed mass and estimated age, the spectroscopic detection of lithium can serve as a marker for DM heating.
Individual systems containing sub-stellar objects that are devoid of hydrogen burning lines are potential sites for detecting dark dwarfs, but an age estimate is also required to distinguish dark dwarfs from young brown dwarfs that have not yet burned their lithium. In practice, a more promising approach would be to apply Bayesian hierarchical modeling e.g., the scheme of [117], to use data from multiple systems to search for a statistical preference for the presence of dark dwarfs.
@@
5 Conclusion
We have found that the structure, evolution, and fate of sub-stellar objects in the presence of dark matter annihilation is markedly different than the Standard Model predictions. Objects lighter than the hydrogen burning limit begin their lives as brown dwarfs but ultimately evolve to become dark dwarfs — eternal objects powered by DM burning. Their radii, temperatures, and luminosity are constant. The H-burning limit itself is modified, depending on the incident DM flux. Stars heavier than this behave similarly to red dwarfs but are predicted to be larger and brighter than SM red dwarfs of identical mass due to the additional contribution from DM heating. Their core temperatures can be reduced, causing them to retain lithium in mass ranges where the SM predicts it is depleted. The detection of lithium-7 in objects heavier than the lithium burning limit would provide evidence for the existence of DM heating. The optimum location to search for these signatures is towards the galactic center, where low DM velocities and large densities maximize the incoming DM flux, with experiments such as JWST (which was shown to have sensitivity to brown dwarfs with temperatures as low as 650 Kelvin [81] in such locations).
@@ Acknowledgements
Appendix A Appendix
In this appendix we review the derivation of the analytic model that we adopted for our calculations above, explore the effects of the assumptions we have made, and derive the range of validity of our model.
A.1 Polytrope Model for the Interior
We begin by reviewing the polytropic description of sub-stellar objects, following reference [88]. The EOS for the combination of electron degeneracy pressure $P_{\rm deg}$ and the gas pressure of the ions $P_{\rm ion}$ is an $n=3/2$ polytrope: [^4] The general polytropic equation of state is given by: $P=K\,\rho^{\frac{n+1}{n}},$ where the polytropic constant $K$ must be calculated from the microphysics.
$P=K\rho^{\frac{5}{3}}\textrm{ with }K=C\mu_{e}^{-5/3}(1+\gamma(\psi)+\alpha\psi).$ (A.1)
@@ A.1 Polytrope Model for the Interior
$C=\frac{2}{5}aA^{\frac{5}{2}},~A=\frac{(3\pi^{2}\hbar^{3}N_{A})^{\frac{2}{3}}}{2m_{e}},~a=\frac{2}{3}\frac{4\pi m_{e}^{\frac{3}{2}}}{(2\pi\hbar)^{3}};$ (A.2)
@@
$\gamma(\psi)=-\frac{5}{16}\psi\ln(1+e^{-\frac{1}{\psi}})+\frac{15}{8}\psi^{2}\left[\frac{\pi^{2}}{3}+\mathrm{Li}_{2}(-e^{-\frac{1}{\psi}})\right]$ (A.4)
with $\mathrm{Li}_{2}$ the polylogarithm of order-2. Using the properties of $n=3/2$ polytropes, it is possible to find expressions for the radius, central temperature, central pressure, and central density as a function of the degeneracy parameter $\psi$ [88]:
$R$ $=R_{0}\mu_{e}^{-5/3}(1+\gamma(\psi)+\alpha\psi)M_{0}^{-\frac{1}{3}}\,$, (A.5)
$T_{c}$ $=T_{0}M_{0}^{\frac{4}{3}}\mu_{e}^{\frac{8}{3}}\frac{\psi}{(1+\gamma(\psi)+\alpha\psi)^{2}}\,$, (A.6)
$\rho_{c}$ $=\rho_{0}M_{0}^{2}\frac{\mu_{e}^{5}}{(1+\gamma(\psi)+\alpha\psi)^{3}}\,$, (A.7)
$P_{c}$ $=P_{0}M_{0}^{\frac{10}{3}}\frac{\mu_{e}^{\frac{20}{3}}}{(1+\gamma(\psi)+\alpha\psi)^{4}}\,$, (A.8)
@@ A.2 Photosphere Models
Table 1: Parameter points for the phase transition [118]. Note that points G and H are very close to the second critical point of hydrogen, where a liquid-liquid phase transition occurs.
The photosphere, which defines the effective temperature and the stellar luminosity, is the point at which the optical depth $\tau$ falls to $2/3$. At this point, there is a phase transition from the mixture of degenerate electrons and ionic gas described above to a phase of molecular hydrogen. Modeling the latter, one can derive expressions for the effective temperature and luminosity. Several parameter points given in table 1 span the possible range of surface temperatures at which the phase transition takes place:
$T_{\rm eff}=b_{1}\times 10^{6}\rho_{e}^{2/5}\psi^{\nu}\rm K,$ (A.9)
where $b_{1}$ and $\nu$ are parameters of the model [118]. Solving the condition $\tau=2/3$ yields an expression for the pressure at the photosphere, which can be used to find the density and effective temperature from the ideal gas law (see [88] for the details – here we generalise the solution presented in this work). The result is:
$T_{\rm eff}=1.57466\times 10^{4}\,b_{1}\psi^{\nu}\left(\frac{M_{0}^{5/3}\psi^{-\nu}}{b_{1}\kappa_{R}(1+\gamma+\alpha\psi)^{2}}\right)^{2/7}\rm K,$ (A.10)
@@
$\frac{L_{\rm surf}}{L_{\odot}}=\frac{0.0578353b_{1}^{3}M_{0}}{\kappa_{R}}\psi^{3\nu}\left(\frac{M_{0}^{5/3}\psi^{-\nu}}{b_{1}\kappa_{R}(1+\gamma+\alpha\psi)^{2}}\right)^{1/7}.$ (A.11)
@@
A.3 Sensitivity to the Molecular Hydrogen Phase Transition
@@
Figure 6: Minimum mass for hydrogen burning (using the criterion $L_{\rm DM}=L_{\rm HB}$) as a function of DM ambient density for different models of the photosphere. Here we use $v_{\rm DM}=$50 km/s.
@@
Our calculation relies upon a choice of parameter point for the phase transition, which introduces a source of theoretical uncertainty into our calculations. We explore this in this section. A convenient measure for quantifying the model-dependency is the minimum mass for hydrogen burning, which is DM density and velocity dependent. Since DM annihilation reduces the core temperature and density below the threshold for PP-burning, the MMHB is increased whenever DM burning is important. Demanding $L_{\rm HB}=L_{\rm DM}$ leads to the minimum mass shown in figure 6 for the different parameter points given in table 1. The figure demonstrates that the model’s predictions only significantly diverge for very large DM densities (with the exception of edge cases G and H), implying that the choice of parameter point is not an important source of uncertainty for our conclusions.
A.4 Coulomb Corrections to the Equation of State
@@
At low temperatures, the effects of Coulomb repulsion becomes important and the EOS deviates from an $n=3/2$ polytrope. This has two consequences for our work. First, it limits the applicability of the analytic model we have assumed to high mass objects and, second, the exponential sensitivity of the lithium-7 burning rate in Eq. (4.2) mandates that we include these corrections to calculate the core temperature accurately. We derive the mass range where our results are valid and expressions for the corrected core temperature in this section.
The effects of Coulomb interactions depend upon the plasma parameter [89]
$\Gamma=\frac{e^{2}}{k_{B}T}\left(\frac{4\pi\rho}{3m_{p}}\right)^{\frac{1}{3}},$ (A.12)
@@
where $e$ is the electron charge and $m_{p}$ is the proton mass. Early in the star’s evolution when lithium is being depleted the plasma parameter $\Gamma\ll 1$, corresponding to the Deybe-Hückel regime [111]. In this limit, each ion can be thought of as being surrounded by a spherically symmetric but inhomogeneously charged cloud that screens its charge. The Coulomb pressure of these screened charges is [108]
$P_{\rm C}=-\frac{e^{2}}{3}\left(\frac{\pi}{k_{B}T}\right)^{\frac{1}{2}}\left(\frac{\rho\zeta}{m_{p}}\right)^{\frac{3}{2}};~\textrm{with }\zeta=\sum_{j}\frac{X_{j}}{A_{j}}Z_{j}(1+Z_{j}),$ (A.13)
@@
where $j$ runs over all ions with mass fraction $X_{j}$, atomic number $A_{j}$ and charge $Z_{j}$$\zeta=1.875$ for our benchmark model. The ratio of the Coulomb pressure to the central pressure calculated using the polytrope model for various DM densities is shown in figure 7. In the SM, the polytropic approximation ceases to be valid ($|P_{\rm C}|>P_{c}$) when $M<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.018}{\rm M}_{\odot}$. This threshold is reduced when DM annihilation is present. The results we presented in this work correspond to heavier objects.
@@
To incorporate the effects of the Coulomb pressure into our lithium calculation, we use that fact that $T_{c}\propto\mu_{\rm eff}^{2}$ [110, 111] with $\mu_{\rm eff}=k_{B}N_{A}\rho_{c}T_{c}/P_{c}$. At fixed central density, the negative Coulomb pressure reduces the total pressure from the polytrope model’s prediction, which raises the central temperature. To incorporate this effect, we first calculate $T_{c}$ using the polytropic model and then scale this by a factor of $P_{c}^{2}/(P_{c}+P_{\rm C})^{2}$ to account for the correction to $\mu_{\rm eff}$.
@@
At late times, when the star has cooled and reached its equilibrium state, either red or dark dwarf, the plasma parameter has evolved to $\Gamma\gg 1$ and the Debye-Hückel approximation does not apply. Instead, the plasma is comprised of a sea of degenerate electrons surrounding a lattice of ions arranged to maximize their separation, minimizing their Coulomb repulsion [89]. In this limit, the Coulomb pressure can be found using the the Wigner-Seitz approximation where each ion with atomic number $Z$ sits at the center of a neutral sphere containing $Z$ electrons. The Coulomb pressure is [89]
@@
$\begin{split}P_{\rm C}&=K_{\rm C}\rho^{\frac{4}{3}};\\ K_{\rm C}&=-\frac{3}{10}\left(\frac{4\pi\langle Z^{2}\rangle}{3}\right)^{\frac{1}{3}}e^{2}\left(\frac{N_{A}}{\mu_{e}}\right)^{\frac{4}{3}}\approx 5.72\times 10^{12}\textrm{g/cm/s}^{2}.\end{split}$ (A.14)
@@
Comparing the central Coulomb pressure to the central pressure (A.8), the latter exceeds the former in objects where $M\lesssim 10^{-3}\,{\rm M}_{\odot}\sim\rm M_{\rm jupiter}$, independent of the DM density. Thus, the equilibrium configurations we have studied are valid at even lower masses than estimated above, the caveat being that the polytrope model provides an invalid description of their formation.
@@
Figure 7: Ratio of the Coulomb to central pressure as a function of mass for varying DM densities. The polytrope model presented in this work becomes invalid when $|P_{\rm C}|/P_{c}>1$.